OpenStudy (anonymous):

find the center of the hyperbola defined by this equation: (y-6)^2/9-(x+8)^2/64=1

OpenStudy (amistre64):

centers are always attached to your xy parts

OpenStudy (amistre64):

(x-Cx) + (y-Cy) = n such that (Cx,Cy) is the center point

OpenStudy (anonymous):

so is that the answer

OpenStudy (amistre64):

im pretty sure it is ...

OpenStudy (amistre64):

what attached to your x part?

hero (hero):

That's not the answer. It's a general form

OpenStudy (anonymous):

i need the answer

OpenStudy (anonymous):

plz

OpenStudy (amistre64):

then tell me what is attached to the x part so we can come to the particular answer ...

OpenStudy (anonymous):

thats all that is given in the question

OpenStudy (amistre64):

yes, so lets read what is given ...

OpenStudy (amistre64):

(x+ ?? ) what do we see attached to the x part?

OpenStudy (anonymous):

i dont no!! whatever is in that question above is all that is given i dont no what x is that is just what is asked thats what i dont understand

OpenStudy (amistre64):

we are not trying to determine what x IS ; but rather what is sitting next to it.

OpenStudy (anonymous):

it wants to no the center of the hyperbola, the answer should be coordinates

OpenStudy (amistre64):

yes, and the coordinates are given in the equation itself; we just have to know where they are sitting in order to simply pluck them from the equation.

OpenStudy (anonymous):

well idk then cuz thats all that is attatched tothis question there is nothing else

OpenStudy (amistre64):

the equation given is this: (y-6)^2/9-(x+8)^2/64=1 lets strip away all the useless noise to get at what we want (y-6) (x+8) now we are left with the parts that we need to get an answer

OpenStudy (amistre64):

what is attached to the x? and what is attached to the y?

OpenStudy (anonymous):

omg idont know

hero (hero):

amistre, by he way, you didn't show him the correct general form.

OpenStudy (anonymous):

i just told you that all was in the question i dont know anything else

OpenStudy (amistre64):

yeah, im sure thats where they are having problems at ...

OpenStudy (amistre64):

you should review your material that shows you how to construct these conic equations ... other than that I dont know what else I can actually do that would be helpful.

hero (hero):

General Form for hyperbola: \[\frac{(y-k)^2}{a^2} - \frac{(x-h)}{b^2} = 1\] where (h,k) is the center In your case, you have: \[\frac{(y-6)^2}{3^2} - \frac{(x-(-8))^2}{8^2} = 1\] where (-8,6) is the center